# Organic or Diffused: Can We Distinguish Human Art from AI-generated Images?

### v2

@@ Abstract.
There are several different approaches to distinguishing human art from AI images, including classifiers trained by supervised learning, research tools targeting diffusion models, and identification by professional artists using their knowledge of artistic techniques. In this paper, we seek to understand how well these approaches can perform against today’s modern generative models in both benign and adversarial settings. We curate real human art across 7 styles, generate matching images from 5 generative models, and apply 8 detectors (5 automated detectors and 3 different human groups including 180 crowdworkers, 4000+ professional artists, and 13 expert artists experienced at detecting AI). Both Hive and expert artists do very well, but make mistakes in different ways (Hive is weaker against adversarial perturbations while Expert artists produce higher false positives). We believe these weaknesses will remain as models continue to evolve, and use our data to demonstrate why a combined team of human and automated detectors provides the best combination of accuracy and robustness.
Figure 1. Samples from curated test set. Human artwork and subsequent matching images produced by generative AI models.
@@ 1. Introduction
All of this begs the question, do we have the tools today to reliably and consistently distinguish AI-generated imagery from human-created art? There are multiple potential solutions. First, human artists are often quite good at recognizing human art, and some experts have demonstrated a consistent ability to detect AI images trying to pass as human art (Gach, 2023; Ong, 2023; Sato, 2023; Shaffi, 2023). Alternatively, specific companies like Hive, Optic and Illuminarty have trained supervised classifiers to distinguish AI imagery from human art. Online media has raised questions on the accuracy of these detectors amid their growing impact on news media (Maiberg, 2023). Finally, recent research results like DIRE and DE-FAKE (Wang et al., 2023a; Sha et al., 2023) suggest specific techniques to recognize images produced by diffusion models, including all of the major generative models today.
@@
In total, we curated a dataset of 280 real human art images across 7 different styles, and 350 generated AI-based images from generative models using prompts automatically extracted from each of the human art images. One component of our study tests the efficacy of automated detectors on these images and their perturbed variants; and the other part evaluates human-based detection. The latter involves 3 separate user studies on: a) 180 crowdworkers on the Prolific platform, b) 4000+ professional artist volunteers recruited from social media artist groups, and c) 13 expert professional artists who have experience identifying generative AI images.
@@ 2.1. Generative Image Diffusion models
First introduced in 2020, diffusion models (Ho et al., 2020) quickly replaced GANs as the state-of-the-art models for image synthesis (Dhariwal and Nichol, 2021). This was quickly followed by extension to text-to-image generation (Ramesh et al., 2022; Nichol et al., 2022). Later instances included multiple open-source models from Stable Diffusion (Rombach et al., 2022; Stability AI, 2022; Podell et al., 2023; StabilityAI, 2022a, b), and commercial models from Midjourney (Midjourney, 2023), DALL-E (OpenAI, 2022, 2023; Betker, 2023), and Adobe Firefly (Adobe, 2023). Diffusion models face a growing number of social and ethical concerns. Base models require enormous amounts of training data, often obtained without consent through web-scraping. Midjourney trained their model using data from over 16,000 artists, the vast majority without consent (Ho, 2024). Stability AI trained on datasets from LAION (Schuhmann et al., 2022; Rombach et al., 2022), containing millions of copyrighted works. These copyright infringement issues have led to multiple class-action lawsuits (CourtListener, 2024; Joseph Saveri Law Firm LLP, 2023). Even in smaller volumes, artists are finding their works being used without consent in finetuned models using techniques such as LoRA (BAIO, 2022). As these models continue to improve in quality, many of their users have attempted to pass AI-generated images as human art. AI-generated images have been used to win art competitions, fooling judges of digital art, photography, and book covers (Roose, 2022; Glynn, 2023; Sato, 2023). Companies that promote human art have found themselves using AI content provided to them by third-party vendors (Orland, 2024).
@@ 2.2. Automated AI-Image Detectors
Research-based detectors.    These come from published research papers that often offer source code and training/testing datasets, sometimes also with pretrained models. While they lack the same public reach and influence as commercial detectors, they offer the benefit of transparency in methodology. Two such detectors are the DIRE detector (Wang et al., 2023a) and DE-FAKE (Sha et al., 2023). DIRE, or Diffusion Reconstruction Error, aims to exploit the forwards and backwards diffusion processes (Ho et al., 2020) to identify images generated by that model (Wang et al., 2023a). DE-FAKE uses both image-only and multimodal embeddings (Radford et al., 2021) to create a model-agnostic detector (Sha et al., 2023).
@@ 2.3. Related work
Detecting Deepfake Photographs.    While our study focuses on distinguishing human art from AI-generated images, several prior studies have focused on human detection of deepfake photos generated by machine learning models. (Bray et al., 2023) evaluates users’ ability to detect deepfakes of human faces using StyleGAN2 (Karras et al., 2020), and finds that human participants have below 65% accuracy in all experiments, even when taught how to recognize deepfakes. Similarly, (Lu et al., 2023) evaluates human detection on a set of real photos and photorealistic images from Midjourney V5 (Midjourney, 2023). They also create a dataset of roughly two million fake images to train ML detectors. While humans misclassify 37% of images, the best-performing ML detector misclassifies 13% of the same evaluation set.
@@ 3.3. Evaluating Human Detection: User Studies
Recent events have shown human artists to be exceptionally successful at identifying AI-generated imagery masquerading as human art (Gach, 2023; Ong, 2023; Sato, 2023; Shaffi, 2023). Instead of looking for statistical properties of images, human artists look for inconsistencies in artistic technique, flaws in logic/composition, and other domain-specific properties that diffusion models do not understand.
@@
  • Baseline Participants. We recruited 180 crowdworkers through the Prolific online crowdsourcing platform (177 completed and passed consistency checks). Participants were compensated \$2/10min and this group took on average 8 minutes to complete. This group included no full-time professional artists and 7 part-time artists.
  • Professional Artist Volunteers. We asked for artist volunteers on social media to participate. Of more than 4000 who responded, 3803 completed the survey and passed all consistency checks.
@@ 4.1. Human Artworks
Human artworks are novel creations by artists that capture their personal touch and emotions. They showcase the unique techniques, styles and perspectives of individual artists that only come from years of training and experiences. With help from the artist community, we collected artworks across 7 major art styles, including anime, cartoon, fantasy, oil/acrylic, photography, sketch, and watercolor. We recruited artist volunteers from Cara (Cara, [n. d.]), a major portfolio platform dedicated for human-made art. When recruiting volunteers, we provided artists with a detailed explanation of the study’s scope and operations. We sought their consent to use their artworks in the study and offered them the option to opt out if they were not comfortable with their works being included.
@@ 4.2. AI-Generated Images
For AI-generated images, we take effort to cover the 7 art styles (listed above) and different AI generators. We consider the five most popular AI generators: CIVITAI (Civitai, 2022), DALL-E 3 (OpenAI, 2023), Adobe Firefly (Adobe, 2023), MidjourneyV6 (Franzen, 2023), and Stable Diffusion XL (SDXL) (Podell et al., 2023). All were the latest release at the time of submission. For each art style, we prompt each AI generator to produce 10 images, for a total of 50 images across all five generators. This produces 350 AI-generated images in total. Table 12 in Appendix summarizes the # of images curated as human artworks and AI-generated images.
@@ 4.3. Perturbed Images
Perturbation #1: JPEG Compression.    Existing work has shown that compression artifacts can reduce the accuracy of image classifiers (Lam et al., 1999; Hamano et al., [n. d.]; Lau et al., 2003). To study its impact on AI image detectors, we follow prior work to apply JPEG compression of a quality factor 15 (Spring et al., 2016) to AI-generated images before querying these classifiers.
@@ 5.2. Results of Unperturbed Imagery
Variation across Automated Detectors.    The large performance variations across detectors could be attributed to the diversity and coverage of their training data. According to its website, Hive utilizes a rich collection of generative AI datasets and can identify the generative model used for the current input from a pool of nine models. Similarly, Optic can pinpoint an input image to one of the 4 generative models. In contrast, Illuminarty’s training data coverage is limited, particularly since it does not support image files exceeding 3MB. In our study, we have to downsample images generated by Firefly from 2048x2048 to 1024x1024 to meet this restriction.
DE-FAKE and DIRE’s training data are constrained and lack representation from art. DIRE models are trained on interior design images (LSUN-Bedroom), human faces (CelebA-HQ) and ImageNet. The detection accuracy on art images is around 50%. For DE-FAKE, we follow the methodology described by (Sha et al., 2023) to train the classifier using images generated by SDXL and MSCOCO captions. When tested on SDXL images produced from MSCOCO prompts, the classifier reproduces a high ACC of 92.44% similar to (Sha et al., 2023). However, when tested on images generated using artwork prompts, the accuracy drops to 50.3% for SDXL images and 46.43% for those produced by other generators.
@@ 5.3. Impact of AI Generator Choice
On the other hand, Firefly images are the least detectable – Hive marks 6 out of 70 Firefly images as human art and 3 as “not sure” while Optic marks 2 as human art and 2 as “not sure” (Fig. 2). We hypothesize this is due to lack of training data. Firefly is a relatively new model, so Hive and Optic are likely to have much less training data relative to other models
@@ 5.4. Impact of Perturbations
Hive 91.88 88.73 94.20 69.73 67.56 96.83
Optic 97.98 52.63 84.21 62.62 58.00 98.85
Illuminarty 93.19 61.43 94.39 89.80 89.22 95.31
Ensemble 94.29 88.86 94.57 73.71 71.43 98.29
@@
Adversarial Perturbation.    Of all the perturbations, the adversarial perturbations on CLIP have the least impact on performance for Hive and Illuminarty. Optic’s ADSR drops to 84.21%, but Hive and Illuminarty both remain above 94%. This may indicate relatively low transferability of attacks in the CLIP space.
@@
Question 2: Does Glaze Affect Accuracy Similarly on Human Artwork vs AI-generated Images    To investigate if the detection of human artwork is impacted similarly to AI-generated images when Glazed, we applied high-intensity Glaze to both. The detection outcomes, including ACC, FPR and FNR are presented in Table 4 comparing results with and without Glaze. Our findings reveal a large reduction in ACC with the use of Glaze, as expected. However, the FNR significantly increases significantly across all detectors, while the FPR remains relatively consistent. On Hive, ACC is decreased around 20% and yet FPR increases only 3.23% while FNR increases almost 30%. Our Ensemble detector is able to achieve the highest accuracy on Glazed images, but maintains a FNR of 28.57%. This suggests that glazing human artwork typically does not affect classification success, but glazing AI-generated images often leads to misclassification as human artwork. We attribute this to the scarcity of Glazed images online and thus the lack of Glazed images in the detector’s training datasets. Additionally Glaze was created to protect human artwork and has since been adopted by many artists, therefore the distribution of Glazed human art vs. Glazed AI-generated images online is likely skewed. Once again this demonstrates the impact of insufficient training data on image-based classifiers.
@@ 6. User Studies on Human Detection
Next, we measure the ability of human users and artists in identifying AI-generated images. As discussed in §3.3, we perform user studies with 3 separate user populations, including crowdworkers, professional artists and expert artists. To ensure uniformity, each group is provided with the same user study, but the expert group received extended followup questions asking for detailed examples (Section 6.4). Our study is approved by our local Institutional Review Board (IRB). We omit the IRB number for anonymity.

### v3

@@ Abstract.
There are several different approaches to distinguishing human art from AI images, including classifiers trained by supervised learning, research tools targeting diffusion models, and identification by professional artists using their knowledge of artistic techniques. In this paper, we seek to understand how well these approaches can perform against today’s modern generative models in both benign and adversarial settings. We curate real human art across 7 styles, generate matching images from 5 generative models, and apply 8 detectors (5 automated detectors and 3 different human groups including 180 crowdworkers, 3800+ professional artists, and 13 expert artists experienced at detecting AI). Both Hive and expert artists do very well, but make mistakes in different ways (Hive is weaker against adversarial perturbations while Expert artists produce higher false positives). We believe these weaknesses will persist, and argue that a combination of human and automated detectors provides the best combination of accuracy and robustness.
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USA.booktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAisbn: 979-8-4007-0636-3/24/10doi: 10.1145/3658644.3670306ccs: Security and privacy Human and societal aspects of security and privacy
Figure 1. Samples of human art and matching images produced by generative AI models. Copyright held by respective artists, ©Kirsty (@kirue_t), ©Nguyen Viet, ©Liam Collod
@@ 1. Introduction
All of this begs the question, do we have the tools today to reliably and consistently distinguish AI-generated imagery from human-created art? There are multiple potential solutions. First, human artists are often quite good at recognizing human art, and some experts have demonstrated a consistent ability to detect AI images trying to pass as human art (Gach, 2023; Ong, 2023; Sato, 2023). Alternatively, specific companies like Hive, Optic and Illuminarty have trained supervised classifiers to distinguish AI imagery from human art. Online media has raised questions on the accuracy of these detectors amid their growing impact on news media (Maiberg, 2023). Finally, recent research results like DIRE and DE-FAKE (Wang et al., 2023a; Sha et al., 2023) suggest specific techniques to recognize images produced by diffusion models, including all of the major generative models today.
@@
In total, we curated a dataset of 280 real human art images across 7 different styles, and 350 generated AI-based images from generative models using prompts automatically extracted from each of the human art images. One component of our study tests the efficacy of automated detectors on these images and their perturbed variants; and the other part evaluates human-based detection. The latter involves 3 separate user studies on: a) 180 crowdworkers on the Prolific platform, b) 3800+ professional artist volunteers recruited from social media artist groups, and c) 13 expert professional artists who have experience identifying generative AI images.
@@ 2.1. Generative Image Diffusion models
First introduced in 2020, diffusion models quickly replaced GANs as the state-of-the-art models for image synthesis (Dhariwal and Nichol, 2021). This was quickly followed by extension to text-to-image generation (Ramesh et al., 2022; Nichol et al., 2022). Later instances included multiple open-source models from Stable Diffusion (Rombach et al., 2022; Stability AI, 2022; Podell et al., 2023; StabilityAI, 2022a, b), and commercial models from Midjourney, DALL-E, and Adobe Firefly.
Diffusion models face a growing number of social and ethical concerns. Base models require enormous amounts of training data, often obtained without consent through web-scraping. Midjourney trained their model using data from over 16,000 artists, the vast majority without consent (Ho, 2024). Stability AI trained on datasets from LAION (Schuhmann et al., 2022; Rombach et al., 2022), containing millions of copyrighted works. These copyright infringement issues have led to multiple class-action lawsuits (CourtListener, 2024; Joseph Saveri Law Firm LLP, 2023). Even in smaller volumes, artists are finding their works being used without consent in finetuned models using techniques such as LoRA (BAIO, 2022).
As these models continue to improve in quality, many of their users have attempted to pass AI-generated images as human art. AI-generated images have been used to win art competitions, fooling judges of digital art, photography, and book covers (Roose, 2022; Glynn, 2023; Sato, 2023). Companies that promote human art have found themselves using AI content provided to them by third-party vendors (Orland, 2024).
@@ 2.2. Automated AI-Image Detectors
Research-based detectors.    These come from published research papers that often offer source code and training/testing datasets, sometimes also with pretrained models. While they lack the same public reach and influence as commercial detectors, they offer the benefit of transparency in methodology. Two such detectors are the DIRE detector (Wang et al., 2023a) and DE-FAKE (Sha et al., 2023). DIRE, or Diffusion Reconstruction Error, aims to exploit the forwards and backwards diffusion processes to identify images generated by that model. DE-FAKE uses both image-only and multimodal embeddings (Radford et al., 2021) to create a model-agnostic detector (Sha et al., 2023).
@@ 2.3. Related work
Detecting Deepfake Photographs.    While our study focuses on distinguishing human art from AI-generated images, several prior studies have focused on human detection of deepfake photos generated by machine learning models. (Bray et al., 2023) evaluates users’ ability to detect deepfakes of human faces using StyleGAN2 (Karras et al., 2020), and finds that human participants have below 65% accuracy in all experiments, even when taught how to recognize deepfakes. Similarly, (Lu et al., 2023) evaluates human detection on a set of real photos and photorealistic images from Midjourney V5. They also create a dataset of roughly two million fake images to train ML detectors. While humans misclassify 37% of images, the best-performing ML detector misclassifies 13% of the same evaluation set.
@@
The most related work on this topic was presented recently at IEEE S&P 2024 (Frank et al., 2024). Where our work focuses entirely on creative visual art, this prior study covered generative AI detection broadly across images, audio and text across Internet users in multiple countries.
@@ 3.3. Evaluating Human Detection: User Studies
Recent events have shown human artists to be exceptionally successful at identifying AI-generated imagery masquerading as human art (Gach, 2023; Ong, 2023; Sato, 2023). Instead of looking for statistical properties of images, human artists look for inconsistencies in artistic technique, flaws in logic/composition, and other domain-specific properties that diffusion models do not understand.
@@
  • Baseline Participants. We recruited 180 crowdworkers through the Prolific online crowdsourcing platform (177 completed and passed attention checks). Participants were compensated \$2/10min and this group took on average 8 minutes to complete. This group included no full-time professional artists and 7 part-time artists.
  • Professional Artist Volunteers. We asked for artist volunteers on social media to participate. Of more than 4000 who responded, 3803 completed the survey and passed all attention checks.
@@ 4.1. Human Artworks
Human artworks are novel creations by artists that capture their personal touch and emotions. They showcase the unique techniques, styles and perspectives of individual artists that only come from years of training and experiences. With help from the artist community, we collected artworks across 7 major art styles, including anime, cartoon, fantasy, oil/acrylic, photography, sketch, and watercolor. We recruited artist volunteers from Cara (Cara, 2023), a major portfolio platform dedicated to human-created art which uses filters to detect AI images and peer-based validation between artists. When recruiting volunteers, we provided artists with a detailed explanation of the study’s scope and operations. We sought their consent to use their artworks in the study and offered them the option to opt out if they were not comfortable with their works being included.
@@ 4.2. AI-Generated Images
For AI-generated images, we take effort to cover the 7 art styles (listed above) and different AI generators. We consider the five most popular AI generators: CIVITAI (Civitai, 2022), DALL-E 3, Adobe Firefly, MidjourneyV6, and Stable Diffusion XL (SDXL). All were the latest release at the time of submission. For each art style, we prompt each AI generator to produce 10 images, for a total of 50 images across all five generators.
@@ 4.3. Perturbed Images
Perturbation #1: JPEG Compression.    Existing work has shown that compression artifacts can reduce the accuracy of image classifiers (Lam et al., 1999; Hamano et al., 2023; Lau et al., 2003). To study its impact on AI image detectors, we follow prior work to apply JPEG compression of a quality factor 15 (Spring et al., 2016) to AI-generated images before querying these classifiers.
@@ 5.2. Results of Unperturbed Imagery
Variation across Automated Detectors.    The large performance variations across detectors could be attributed to the diversity and coverage of their training data. According to its website, Hive utilizes a rich collection of generative AI datasets and can identify the generative model used for the current input from a pool of nine models  [^2] As of April when conducting additional tests on Hive, it now detects 27 models.. Similarly, Optic can pinpoint an input image to one of the 4 generative models. In contrast, Illuminarty’s training data coverage is limited, particularly since it does not support image files exceeding 3MB. In our study, we have to downsample images generated by Firefly from 2048x2048 to 1024x1024 to meet this restriction. Illuminarty and Optic are smaller companies with less training data  [^3] Illuminarty’s creator informed us the model has not been updated for 6 months in Novemeber 2023..
DE-FAKE and DIRE’s training data are constrained and lack representation from art. DIRE models are trained on interior design images (LSUN-Bedroom), human faces (CelebA-HQ) and ImageNet. The detection accuracy on art images is around 50%. For DE-FAKE, we follow the methodology described by (Sha et al., 2023) to train the classifier using images generated by SDXL and MSCOCO captions. When tested on SDXL images produced from MSCOCO prompts, the classifier reproduces a high ACC of 92.44% similar to (Sha et al., 2023). However, when tested on images generated using artwork prompts, the accuracy drops to 50.3% for SDXL images and 46.43% for those produced by other generators. We attribute performance discrepancy to the issue of transferability. Training images generated from MSCOCO prompts do not follow the art dataset distribution, so these open-source detectors did not transfer well to the art dataset.
@@ 5.3. Impact of AI Generator Choice
On the other hand, Firefly images are the least detectable – Hive marks 6 out of 70 Firefly images as human art and 3 as “not sure” while Optic marks 2 as human art and 2 as “not sure” (Fig. 2).
We hypothesize this is due to lack of training data. Firefly is a relatively new model, so Hive and Optic are likely to have much less training data relative to other models
@@ 5.4. Impact of Perturbations
Hive 91.88 88.73 93.00 69.73 67.56 96.83
Optic 97.98 52.63 80.42 62.62 58.00 98.85
Illuminarty 93.19 61.43 94.34 89.80 89.22 95.31
Ensemble 94.29 88.86 94.29 73.71 71.43 98.29
@@
Adversarial Perturbation.    Of all the perturbations, the adversarial perturbations on CLIP have the least impact on performance for Hive and Illuminarty. Optic’s ADSR drops to 80.42%, but Hive and Illuminarty both remain above 93%. This may indicate relatively low transferability of targeted attacks in the CLIP space in black-box settings.
@@
Question 2: Does Glaze Affect Accuracy Similarly on Human Artwork vs AI-generated Images?    To investigate if the detection of human artwork is impacted similarly to AI-generated images when Glazed, we applied high-intensity Glaze to both. The detection outcomes, including ACC, FPR and FNR are presented in Table 4 comparing results with and without Glaze. Our findings reveal a large reduction in ACC with the use of Glaze, as expected. However, the FNR increases across all detectors, while the FPR remains relatively consistent. On Hive, ACC is decreased around 20% and yet FPR increases only 3.23% while FNR increases almost 30%. Our Ensemble detector is able to achieve the highest accuracy on Glazed images, but maintains a FNR of 28.57%. This suggests that glazing human artwork typically does not affect classification success, but glazing AI-generated images often leads to misclassification as human artwork. We attribute this to the scarcity of Glazed images online and thus the lack of Glazed images in the detector’s training datasets. Additionally Glaze was created to protect human artwork and has since been adopted by many artists, therefore the distribution of Glazed human art vs. Glazed AI-generated images online is likely skewed. Once again this demonstrates the impact of insufficient training data on image-based classifiers.
@@ 5.5. Summary of Findings
Figure 4. CDF of overlay intensity required to change Hive’s decision over the period of 3 months.
Figure 5. CDF of overlay intensity required to change Hive’s decision using different overlay methods in June 2024.
Figure 6. User study interface.
5.6. Followup Robustness Tests on Hive
After the paper was accepted, we discovered a Reddit post that claimed overlaying a real image of a white wall onto an AI-generated image could bypass Hive’s detection (YentaMagenta, 2023). We investigate this as a new adversarial approach to bypassing detection, and conduct a large-scale test on 105 AI-generated images, randomly sampled from our dataset to include 3 images from each of the 7 styles and 5 AI-generators. We implement a python script to overlay each image with a white wall image from Adobe Shutterstock, at intensity levels varying from 10% to 90% in increments of 10. Hive confirms the Adobe’s white wall image as not AI-generated with 100% confidence.
Figure 6 illustrates the CDF graph of the distribution of overlay intensities required to flip Hive’s decision from AI-generated to not AI-generated. The red curve from April 2024 shows there is a 17.5% chance that Hive’s decision will change at or below intensity levels of 50%. Yet, there are corner cases where images (2.91% in April and 3.85% in June) were able to fool Hive with the white wall overlaid at only 10% intensity. Most images require around overlay intensity of 60-80% to change Hive’s classification.
We also note that Hive’s performance improved over time. Figure 6 displays Hive results collected in April 2024 and June 2024. The shift in the CDF shows that Hive’s accuracy and robustness increased in the 2 month period, possibly due to adversarial training on inputs or updates to its supervised classification algorithm.
We conduct another smaller-scale test to see the effect of different overlay methods on Hive’s output. We apply and compare Adobe Photoshop’s “multiply blend” overlay option (the method used in the Reddit post) against our scripted blending algorithm, varying intensities from 10% to 40% in increments of 10 across 63 AI-generated images [^4] Dataset is narrowed down to 3 images from MJv6, Firefly, and SDXL across 7 styles.. Figure 6 shows Adobe’s Photoshop method has a significantly stronger effect against Hive at higher intensities. This gap is somewhat unexpected, and just confirms that even small variations in adversarial attacks can have large unpredictable impact on Hive’s detection robustness.
@@ 6. User Studies on Human Detection
Next, we measure the ability of human users and artists in identifying AI-generated images. As discussed in §3.3, we perform user studies with 3 separate user populations, including crowdworkers, professional artists and expert artists.
To ensure uniformity, each group is provided with the same user study, but the expert group received extended followup questions asking for detailed examples (Section 6.4). Our study is approved by our local Institutional Review Board (IRB). We omit the IRB number for anonymity.